# Mean field model of contagion processes in urban traffic networks

### v1

@@ Abstract
Theoretical arguments and empirical evidence for the emergence of macroscopic epidemic type behavior, in the form of Susceptible-Infected-Susceptible (SIS) or Susceptible-Infected-Recovered (SIR) processes in urban traffic congestion from microscopic network flows is given. Moreover, it’s shown that the emergence of SIS/SIR implies a relationship between traffic flow and density, which is consistent with observations of the so called Fundamental Diagram of Traffic, which is a characteristic signature of vehicle movement phenomena that spans multiple scales. Our results provide a plausible explanation for this scale-spanning signature and put in more firm grounds recent findings that indicate that traffic congestion at the aggregate level can be modeled by simple contagion dynamics.
@@ 1 Introduction
Vehicular movement in urban networks can be understood like an out of equilibrium many-body system with a hierarchy of description levels and relevant scales. From a microscopic standpoint, traffic is a granular flow with strong interactions among particles [1]. These interactions are in part a consequence of decision making of the individual particles themselves, which in this sense can be regarded as agents whom can display non-linear and stochastic behavior. At a macroscopic level on the other hand, traffic can be regarded as a continuous flow in terms of coarse grained observables like densities and average velocities, being perhaps manageable by means of interventions on road geometry and network topology intended to reduce traffic jams [2, 3, 4]. Urban traffic capacity is characterized by a Fundamental Diagram of Traffic (FDT), which indicates the existence of a free-flow phase and a congested phase with a transition between them in a critical vehicle density point. This characteristic FDT has been observed to span several scales ranging from the individual street level up to aggregate dynamics over the urban network. At the largest urban scales, traffic congestion has been shown to display a remarkable empirical correspondence with epidemics and propagation by contagion processes [5, 6, 7]. In the present work we put forward a possible unifying framework that explains the scale spanning FDT and the epidemic large scale dynamics by identifying them as emergent properties of an out of equilibrium system that shares commonalities with both granular and continuous flows. The vicinity of road intersections are unavoidable sources of discontinuity and therefore, the places at which the granular nature of the flow becomes more apparent. In other regions far away from the intersections, the continuous density description dominates if additional sources of instability, like bottlenecks, are absent. Therefore we propose a microscopic contact at the crossroads mechanism for the transmission of flow in a transport network. Our resulting model, detailed in the next section, is conceptually similar to the point queue and traffic flow models (see for instance [8] for a discussion on these approaches) with the crucial difference that the transmission rate parameters are not defined by local characteristics of the intersection. In our contagion at junctions model, the rate parameters are instead determined by the total transport in a given time period and the average speed in the road that receives the vehicles. The rate parameters are therefore a consequence of the local discontinuity produced at junctions where incoming vehicles must adapt to the average flow of the receiving road, being this average flow determined by non local spatial and temporal constraints.
@@
2 Crossroad contagion network traffic flow model
@@
Urban traffic has been shown to be consistent with an out of equilibrium granular flow process [1]. Consider therefore a network of links (edges), each representing idealized roads through which such a flow occurs. It’s additionally assumed that aggregates of cars traversing a link $n$ can be described in terms of a continuous car density flow at any point of the link except at crossroads, which can be regarded as contact points in which vehicles enter or leave the $n$-th link. The master equation, or instantaneous transition time limit of the Chapman-Kolmogorov equation [9], for the car density at an edge $n$ is given by,
@@
The $q_{t}(r|l)$ terms in Eq.(1) represent the instantaneous flow rate of cars from an edge $l$ to another directly connected edge $r$. We now show how the general flow traffic formulation given by Eq.(1), leads naturally to a contagion type model. Consider an intersection in which a link $l$ sends its vehicles to a directly connected link $r$. Assume that there is no other link connected to $r$ besides $l$ in a neighborhood at a distance $\delta x_{r}$ from the intersection $(r,l)$ (see the Figure 1). Denote by $\beta_{r}\equiv\frac{\delta x_{r}}{\delta t}$, the instantaneous average car speed at the interval $\delta x_{r}$ of the $r$-th edge. The number of vehicles $\delta c_{r,l}(t)$ that enter $r$ from $l$ during a time interval $(t,t+\delta t)$ can be expressed equivalently in terms of the flux from $l$ to $r$, denoted by $q(r|l)$, or in terms of the average transiting velocity in the intersection neighborhood $\delta x_{r}$,
$\delta c_{r,l}(t)=q(r|l)\delta t=\frac{q(r|l)\delta x_{r}}{\beta_{r}}$ (2)
Therefore,
$q_{t}(r|l)=\beta_{r}\frac{\delta c_{r,l}}{\delta x_{r}}$ (3)
Now let’s introduce the additional assumption that at the vicinity of the intersection $(r,l)$, the vehicle density of the $r$-th link, $\rho_{r}$ is given by $\rho_{r}=\frac{\delta c_{r,l}}{\delta x_{r}}$. This assumption simply reflects the model setup in which there is no vehicle source for the road $r$ other than road $l$ in the neighborhood $\delta x_{r}$. Equation (3) then leads to the relation,
$q_{t}(r|l)=\beta_{r}\rho_{r}$ (4)
From equation (4), the vehicle density increment $\delta\rho_{r}$ in the time interval $\delta t$, in the road $r$ due to the incoming flow from the intersecting road $l$ is given by
$\delta\rho_{r}=\beta_{r}\rho_{r}\rho_{l}\delta t$ (5)
Now suppose that the $r$-th edge has another crossroad by which it’s directly connected to a link $k$ that acts like the venting point for its flow. By the same arguments that lead to the equation (5), the vehicle density change in the road $r$ due to its intersections with $l$ and $k$ is now given by,
$\frac{\delta\rho_{r}}{\delta t}=\beta_{r}\rho_{r}\rho_{l}-\beta_{k}\rho_{k}\rho_{r}$ (6)
By the equation (6), the master equation (1) can be written in terms of the adjacency matrix of directed links that send flow towards a given link $i$, $\mathcal{A}(i|j)$ as,
$\frac{d\rho_{n}(t)}{dt}=\sum_{m\neq n}\mathcal{A}(n|m)\beta_{n}\rho_{n}\rho_{m}-\sum_{k\neq n}\mathcal{A}(k|n)\beta_{k}\rho_{k}\rho_{n}$ (7)
$=\rho_{n}\left(\beta_{n}\sum_{m\neq n}\mathcal{A}(n|m)\rho_{m}-\sum_{k\neq n}\mathcal{A}(k|n)\beta_{k}\rho_{k}\right)$
Note that the second term inside the parenthesis of Eq.(7) acts like a venting rate for the edge $n$, call it $\gamma_{n}$. Regarding the first term inside of the parenthesis, it can be rewritten by the use of the identity $\sum_{m}\mathcal{A}(n|m)\rho_{m}=\rho_{n}+\sum_{m\neq n}\mathcal{A}(n|m)\rho_{m}$. By the given considerations, the equation (7) therefore reads,
$\frac{d\rho_{n}(t)}{dt}=\rho_{n}\left[\beta_{n}\left(\langle\rho(t)\rangle-\rho_{n}\right)-\gamma_{n}\right]$ (8)
In the previous equation, the quantity $\langle\rho(t)\rangle$ represents the average density of vehicles inside the sub graph formed by $M$ directed links that send vehicles to $n$ together with the $n$-th link itself,
$\langle\rho(t)\rangle\equiv\sum_{m=1}^{M}\mathcal{A}(n|m)\rho_{m}(t)\equiv\bar{\rho}$ (9)
The equation 9 should be interpreted as the average density traversing the link $n$ arrived from its directly connected roads in a time interval $\delta t$. Under an homogeneous mean field description,
$\lim_{M\to\infty}\sum_{m=1}^{M}\mathcal{A}(n|m)\rho_{m}(t)=\bar{\rho}$ (10)
The homogeneous mean field vehicle density dynamics on the $n$-th road is therefore governed by,
$\frac{d\rho_{n}}{dt}=\rho_{n}\left[\beta_{n}(\bar{\rho}-\rho_{n})-\gamma_{n}\right]$ (11)
[…]
@@
[…]
$\rho_{n}(\infty)=0,\quad 0\leq\rho_{n}(t)\leq 1\quad\forall\quad n,t.$
The constraints in Eq. (12) model the distinctiveness of each $24$ hour time span due to the weekly seasonality. These constraints can be satisfied by introducing additional variables defined through,
$I\equiv N_{n}-S-R,$ (13)
[…]
In the above equations, $N_{n}$ is the total transport on link $n$ during a given $24$ hour time span. The definitions given by Eq.(13) imply that,
$\frac{dS}{dt}+\frac{dI}{dt}+\frac{dR}{dt}=0,$ (14)
$s_{n}(t)+\rho_{n}(t)+r_{n}(t)=C.$
By choosing $C=1$ without loss of generality, the constrained Eq. (12) directly leads to a Susceptible-Infected-Recovered (SIR) type model for the traffic flow over the $n$-th node, in which each daily shift can be interpreted as an individual contagion process,
$\dot{s}_{n}(t)$ $=$ $-\beta_{n}\rho_{n}(t)s_{n}(t)$ (15)
$\noindent\dot{\rho}_{n}(t)$ $=$ $\beta_{n}\rho_{n}(t)s_{n}(t)-\gamma_{n}\rho_{n}(t)$
$\noindent\dot{r}_{n}(t)$ $=$ $\gamma_{n}\rho_{n}(t).$
In the above equations, $s_{n}$ can be interpreted as the instantaneous density of vehicles in the sub graph conformed by the incoming directly connected links to $n$ and $r_{n}$ the instantaneous density of vehicles in the sub graph conformed by the directly connected links outwards from $n$. The constant flow setup equation (11), is on the other hand equivalent to a SIS contact process in which $\rho_{n}$ is interpreted as the probability that the edge $n$ becomes infected and $1-\rho_{n}$ is the probability of the edge $n$ being susceptible. The SIS/SIR descriptions can be expected to hold in a macroscopic limit under vanishing fluctuations. Define $\beta_{M}=\frac{1}{M}\sum_{m=1}^{M}\beta_{m}$, the average transiting speed over an entire network with $M$ roads. Define also, $\gamma_{M}=\frac{1}{M}\sum_{m=1}^{M}\gamma_{m}$. The average of the dynamics over the graph $\frac{d\rho_{M}}{dt}=\frac{1}{M}\sum_{m}\frac{d\rho_{m}}{dt}$, then reads,
$s_{M}(t)$ $=$ $s_{M}(0)e^{-\beta_{M}\rho_{M}(t)}$ (16)
$\frac{d\rho_{M}(t)}{dt}$ $=$ $\beta_{M}\rho_{M}(t)s_{M}(t)-\gamma_{M}\rho_{M}(t)+\left[\left\langle\left(\frac{d\rho_{M}}{dt}\right)^{2}\right\rangle-\left\langle\left(\frac{d\rho_{M}}{dt}\right)\right\rangle^{2}\right]$
Under vanishing fluctuations at large $M$, the average dynamics over the network converges to the classical compartmental SIR process.
Figure 1: Relation for the the total number of vehicles traversing the $n$-th link with respect to the instantaneous share of vehicles given by its incoming connected roads and expelled to its outwardly directed roads. The contagion at the crossroads model for traffic proposes that the incoming vehicles should adapt to the average speed of the receiving link, which make the crossroads contact points lasting sources of instability.
@@ 2.1 Comparison with related approaches and the Fundamental Diagram of Traffic
Descriptions of traffic based on contagion dynamics models have been previously proposed, but these depend on ad-hoc definitions of congested states. For instance, in the seminal work [5], congestion is defined by mapping to bond percolation on a small world network, where the congested states can be transmitted according to a contagion rate parameter and a percolation probability parameter. More recently in [6], congested states are formalized in terms of a threshold for the instantaneous average speed on a link relative to its maximum speed. Given this definition for a congested state road, is then proposed in [6] that congestion spreads through the network by a mechanism that at the aggregated level is described by a simple compartmental SIR dynamics. Ample empirical evidence is given in [6], that indicates that the compartmental SIR behavior closely match the temporal evolution of congestion in a given city.
@@
The resulting shape is consistent with the Fundamental Diagram of Traffic (FDT) [8], in the sense that it permits to distinguish between free flow and congested phases by the sign of the derivative of the flow with respect to the density. See the upper panel of the Figure 2 for an empirical estimation of an FDT at a local single detector mean field level.

### v3

# Contagion mean field model for transport in urban traffic networks

@@ Abstract
Theoretical arguments and empirical evidence for the emergence of macroscopic epidemic type behavior, in the form of Susceptible-Infected-Susceptible (SIS) or Susceptible-Infected-Recovered (SIR) processes in urban traffic congestion from microscopic network flows is given. Moreover, it’s shown that the emergence of SIS/SIR implies a relationship between traffic flow and density, which is consistent with observations of the so called Fundamental Diagram of Traffic (FDT), which is a characteristic signature of vehicle movement phenomena that spans multiple scales. Our results put in more firm grounds recent findings that indicate that traffic congestion at the aggregate level can be modeled by simple contagion dynamics.
@@ 1 Introduction
Vehicular movement in urban networks can be understood like an out of equilibrium many-body system with a hierarchy of description levels and relevant scales. From a microscopic standpoint, traffic is a granular flow with strong interactions among particles [nobottle]. These interactions are in part a consequence of decision making of the individual particles themselves, which in this sense can be regarded as agents which can display non-linear and stochastic behavior. At a macroscopic level on the other hand, traffic can be regarded as a continuous flow in terms of coarse grained observables like densities and average velocities, being perhaps manageable by means of interventions on road geometry and network topology intended to reduce traffic jams [cassidy, dganzo, loder]. Urban traffic capacity is characterized by a Fundamental Diagram of Traffic (FDT), which indicates the existence of a free-flow phase and a congested phase with a transition between them in a critical vehicle density point, as further illustrated in the simple triangular version of the flow vs density FDT schematics provided in the Figure 1 This characteristic FDT has been observed to span several scales ranging from the individual street level up to aggregate dynamics over the urban network. At the largest urban scales, traffic congestion has been shown to display a remarkable empirical correspondence with epidemics and contagion propagation processes, by fitting observed congestion to compartmental epidemic models [saberi, duan]. In the present work we put forward a possible unifying framework that explains the scale spanning FDT and the epidemic large scale dynamics by identifying them as emergent properties of an out of equilibrium system that shares commonalities with both granular and continuous flows. The vicinity of road intersections are unavoidable sources of discontinuity and therefore, the places at which the granular nature of the flow becomes more apparent. In other regions far away from the intersections, the continuous density description dominates if additional sources of instability, like bottlenecks, are absent. Therefore we propose a microscopic contact at the crossroads mechanism for the transmission of flow in a transport network. Our resulting model, detailed in the next section, is conceptually similar to the point queue and traffic flow models (see for instance [blubook-vol1-v10] for a discussion on these approaches) with the crucial difference that the transmission rate parameters are not defined by local characteristics of roads and junctions. Moreover, in contrast to previous approaches that use epidemic models as a metaphor for traffic modeling [sun, saberi, duan], our proposed formalism introduced in Section 2, shows that the mathematical conservation of flux at intersections naturally yields the FDT shape. Empirical evidence for this result is provided in Section 3. Our first principles derivation of the Fundamental Diagram of Traffic has theoretical interest on its own and also potential applications for traffic control, as discussed in Sections 2 and 4. By identifying traffic congestion as mathematically equivalent to a transmitted disease, our model gives possible immunization strategies.
Figure 1: Schematic representation of the Fundamental Diagram of Traffic (FDT). The triangular shape illustrates the two distinct regimes: the free-flow phase (blue solid line) where flow increases linearly with density, and the congested phase (red dashed line) where interactions reduce flow as density approaches the jam density ($\rho_{jam}$). The peak represents the road capacity ($q_{max}$) at the critical density ($\rho_{c}$).
@@
2 Mean field network traffic flow model of contagion at the crossroads
Figure 2: Junctions level illustration of the contagion traffic flow model. In the mean field limit, the aggregation of inflow and outflow intersections like those in the Figure, give a coarse grained description equivalent to a Susceptible-Infected-Susceptible (SIS) contagion process under stationary flow conditions. If the total transport is constrained to occur in a given time interval, the mean field model is then equivalent to a Susceptible-Infected-Recovered (SIR) process.
@@ 2 Mean field network traffic flow model of contagion at the crossroads
Urban traffic has been shown to be consistent with an out of equilibrium granular flow process [nobottle]. Consider therefore a network of links (edges), each representing idealized roads through which such a flow occurs. It’s additionally assumed that aggregates of cars traversing a link $n$ can be described in terms of a continuous car density flow in any point of the link except at the crossroads, which can be regarded as contact points in which vehicles enter or leave the $n$-th link. The master equation, or instantaneous transition time limit of the Chapman-Kolmogorov equation [VanKampen], for the car density at an edge $n$ is given by,
@@
The $q_{t}(n|m)$ and $q_{t}(r|n)$ terms in Eq.(1) represent the instantaneous flow rates of cars from the edges $m$ and $n$ to their directly connected edges $n$ and $r$, respectively. The precise description of $q_{t}(n|m)$ and $q_{t}(r|n)$ involve nonlinear interactions of vehicles at both sides of a given juncture due to the compromise between the vehicles in both lanes. In this sense, traffic signaling can be understood like a regulator of the inherent instabilities at street crossroads. Consider the flows, densities and average speeds in the vicinity of the intersections $n\leftarrow m$, where the arrow indicates transport from $m$ to $n$, $m\neq n$ and at $r\leftarrow n$, $r\neq n$ which vent transport outside the $n$-th road. In the present work we propose an average description of the traffic network junctions such that pockets of vehicles with densities $\rho_{k\leftarrow u}$ are transmitted in an intersection $k\leftarrow u$ with an average rate of $\bar{v}_{k\leftarrow u}$, so
$q_{t}(n|m)=\bar{v}_{n\leftarrow m}\rho_{n\leftarrow m}(t)$ (2)
$q_{t}(r|n)=\bar{v}_{r\leftarrow n}\rho_{r\leftarrow n}(t)$
The Eq. (2) can be interpreted like an extension of the definitional relationship between flow and density [libre, blubook-vol1-v10], $q=\bar{v}\rho$. Our proposed description broadens the definitional relationship by considering the $\bar{v}_{k\leftarrow u}$ terms to be the transport rates of densely packaged groups of vehicles at the neighborhood of a given intersection $k\leftarrow u$. Substituting the Eq. (2) in the master equation Eq. (1),
$\frac{d\langle\rho_{n}\rangle}{dt}=\sum_{m\neq n}\langle\bar{v}_{n\leftarrow m}\rho_{n\leftarrow m}\rho_{m}\rangle-\sum_{r\neq n}\langle\bar{v}_{r\leftarrow n}\rho_{r\leftarrow n}\rho_{n}\rangle$ (3)
The brackets in Eq. (3) denote spatial averages over the joint density of vehicles traversing the network. In the zeroth order approximation of vanishing vehicle density fluctuations, the mean field dynamics reads,
$\frac{d\langle\rho_{n}\rangle}{dt}=\langle\rho_{n}\rangle\left[\sum_{m\neq n}\langle\bar{v}_{n\leftarrow m}\rangle\langle\rho_{m}\rangle-\sum_{r\neq n}\langle\bar{v}_{r\leftarrow n}\rangle\langle\rho_{r}\rangle\right]$ (4)
The last term inside the square brackets in Eq. (4) is simply the venting rate of vehicles outside the $n$-th link. The first term inside the brackets gives the total vehicle transmission of incoming vehicles from links other than $n$. By normalizing spatially over the subgraph of links directly connected to $n$, the following coarse grained dynamics is obtained,
$\frac{d\rho_{n}(t)}{dt}=\beta_{n}\rho_{n}(1-\rho_{n})-\gamma_{n}\rho_{n}$ (5)
The mean field expression Eq. (5) is equivalent to a Susceptible-Infected-Susceptible (SIS) contagion process by identifying $\beta_{n}$ and $\gamma_{n}$ like contagion and recovery rate parameters, respectively. In the present context, these refer to the effective transmission and venting of vehicles to a link $n$ by its connected network of inwards and outwards links. If a total number of vehicles $C_{n}$ traverse the $n$-th link during a time period $T$, then under conservative conditions the density $\rho_{n}(t)$ should be normalized in time as,
@@
$\int_{0}^{T}\rho_{n}(t)dt=C_{n}$
The constraints in Eq. (6) can be satisfied by introducing the temporal normalization factors,
$I_{n}\equiv\int_{0}^{T}\rho_{n}(t)dt,$ (7)
$S_{n}\equiv\int_{0}^{T}(C_{n}-\rho_{n})dt\equiv\int_{0}^{T}s_{n}(t)dt$
$R_{n}\equiv C_{n}-S_{n}-I_{n}$
The Eq. (7) states that the transport that traverse the $n$-th link during a time interval $T$ is conserved. The introduction of the constraints Eq.(6) therefore lead to a Susceptible-Infected-Recovered (SIR) type model for the traffic flow over the $n$-th node, in which a time span $T$ with total transport $C_{n}$ can be interpreted as an individual contagion process. By combining with the vanishing spatial density fluctuations limit of Eq. (5), the following coarse grained SIR description of an individual link in a mean field network is obtained,
$s(t)$ $=$ $s(0)e^{-\beta\rho(t)}$ (8)
$\frac{d\rho(t)}{dt}$ $=$ $\beta\rho(t)s(t)-\gamma\rho(t)$
@@ 2.1 Comparison with related approaches and the Fundamental Diagram of Traffic
Descriptions of traffic based on contagion dynamics models have been previously proposed, but these depend on ad-hoc definitions of congested states. For instance, in the seminal work [sun], congestion is defined by mapping to bond percolation on a small world network, where the congested states can be transmitted according to a contagion rate parameter and a percolation probability parameter. More recently in [saberi], congested states are formalized in terms of a threshold for the instantaneous average speed on a link relative to its maximum speed. Given this definition for a congested state road, is then proposed in [saberi] that congestion spreads through the network by a mechanism that at the aggregated level is described by a simple compartmental SIR dynamics. Ample empirical evidence is given in [saberi], that indicates that the compartmental SIR behavior closely match the temporal evolution of congestion in a given city.
@@
The resulting shape is consistent with the Fundamental Diagram of Traffic (FDT) [blubook-vol1-v10], in the sense that it permits to distinguish between free flow and congested phases by the sign of the derivative of the flow with respect to the density.
@@
The result Eq. (10) clearly expresses the value of our mean field approach for the understanding of traffic in terms of contagion dynamics. Previous epidemic models of road networks, like the one presented in [saberi], are largely reactive: they track the spread of a jam after certain speed threshold is breached. Our result Eq. (10) in contrast, offers a predictive capability. It implies that the “tipping point” into congestion is not a universal constant but is a function of the inflow or infection rate ($\beta$), the outflow or recovery rate ($\gamma$) and $s_{0}$ (the network load). This suggests that a traffic controller could theoretically increase the critical density (making the road more resistant to jams) by altering $\gamma$ (e.g., adjusting traffic light cycles to increase venting) relative to $\beta$. This provides a theoretical basis for “immunizing” intersections against jams by tuning their topological parameters, rather than just managing queues once they form.
@@ 3 Comparisons with empirical data
3.1 Direct experimental evidence
To have data for the direct empirical evidence of the proposed mean field model at the single link level requires precise flow measurements from all the relevant intersections in an entire subgraph. This can be difficult due to the usually incomplete coverage provided by traffic sensors. Here however we present an empirical study that intends to overcome the stated obstacle by analyzing vehicle flow data in a lane segment close to an intersection where a group of important traffic links meet. The case study is situated in the city of Saltillo, capital of the state of Coahuila in Mexico. The data has been gently provided by Autonomous University of Coahuila professor Dr. Jaime Burgos García together with the Municipal Institute of Sustainable Urban Mobility of Saltillo. The experimental setup has been implemented as follows. A vehicle count station for the road segment of interest, which is part of the Venustiano Carranza Boulevard in Saltillo, has been situated at the approximate latitude and longitude coordinates (25.459628,100.983199)(25.459628,-100.983199). Another vehicle count station has been placed nearly $1.0$ Km in direction north, at the approximate coordinates (25.466904,100.979632)(25.466904,-100.979632). The second station monitors the already mentioned hub of lanes, one of those corresponding to the chosen road section of the Venustiano Carranza Boulevard. By the combined vehicle count measurements of both stations, which haven taken in $15$ minutes time intervals, the number of cars entering and leaving the road segment through the hub and the total number of vehicles in the segment have been registered. The aggregated data is presented in the Table 1, in which the number of vehicles traversing the lane segment at the hub intersection is calculated as the difference between the total input plus output vehicles and the number of vehicles incoming from the Venustiano Carranza Boulevard road segment itself and continuing their transit along it. A total of $20$ measurements have been made the $25$th of September, $2024$. The first ten measurements have been taken at the morning traffic peak, which goes from $7:00$ AM to $9:30$ AM, while the second half of the dataset instances correspond to the afternoon peak hour, from $17:00$ PM to $19:30$ PM, local times.
We have used the provided data to test our mean field contagion model by normalizing the vehicles counts per measurement considering the total transport through the intersection during the total observation time span. Therefore the car density flux at the intersection is normalized as,
$\Delta\rho/\Delta t=\frac{(\text{number of vehicles traversing the intersection})}{(20*\text{total transport})}$ (11)
In this way, the data in the last column of the Table 1 gives a vehicle flux time series that we have used to fit the discrete Euler approximation of the SIR model, Eq.(8), neglecting fluctuations. The results are shown in the Figure 3, where the dotted lines represent the flow data, the dashed blue lines its 9090% confidence interval and the middle dashed red line gives the fitted SIR model. The predicted average rate parameters turn out to be $\beta=24.435\pm 4.0$, $\gamma=23.205\pm 4.0$, which are consistent with the actual observed rates of $25$ and $27$ vehicles per minute, respectively. Figure 4 on the other hand, uses the predicted mean $\beta$ and $\gamma$ rates to evaluate the function $\dot{\rho}(\rho)$, which gives the Fundamental Diagram of Traffic (FDT) for the considered road segment at the considered time spans. The large confidence intervals shown in Figure 3 confirm that significant fluctuations are present, which naturally leads to scatter in the FDT representation of Figure 4. Besides the statistical noise resulting from the limited sample size, the empirical data also displays the phenomenon of hysteresis loop in the Fundamental Diagram, which is considered to be a consequence of drivers behavior during acceleration and deceleration [traffhys]. It’s interesting to notice that our model successfully identifies the attractor (the red line) around which the real, noisy, hysteretic data (the black line) oscillate. That is, our mean field model successfully captures the central tendency (the “backbone”) of the Fundamental Diagram, effectively averaging through microscopic interactions like those resulting in the hysteresis loop of the experimental data.
Table 1: Vehicles count data at the peak hours in the vicinity of a highways intersection in Saltillo, Mexico
15 minutes intervals Input vehicles Output vehicles Vehicles in the Vehicles traversing
segment of link the intersection
1 466 482 418 530
2 502 461 422 541
3 444 435 372 507
4 521 486 427 580
5 445 413 344 514
6 428 422 359 491
7 364 384 307 441
8 454 424 363 515
9 373 373 317 429
10 256 295 223 328
11 294 369 228 435
12 373 461 275 559
13 356 466 270 552
14 334 446 268 512
15 253 370 182 441
16 221 303 162 362
17 204 218 131 291
18 284 323 230 377
19 343 415 295 463
20 443 466 353 556
Average rates per minute 25 27
Total transport 9424
Figure 3: Fitting of the SIR mean field model to experimental data in the city of Saltillo, Mexico (see Subsection 3.1 for a thorough explanation).
Figure 4: Experimental vehicle flow vs density data from the city of Saltillo, Mexico (black dotted lines) and the FDT function that results from the estimated SIR model (red dotted dashed lines).
3.2 Comparisons with traffic sensors data